OpenStudy (anonymous):

For the pair of functions, f and g, find (f o g)(x) and (gof)(x). f(x) =4-7x,g(x)=5x-5 Please help!Thanks in advance.

OpenStudy (neer2890):

(f o g)(x)= f(g(x)) so in f(x) put g(x) in place of x. for e.g., (f o g)(x)= f(g(x))=f(5x-5)=4-7(5x-5) so (f o g)(x) = 4-35x+35=39-35x

OpenStudy (neer2890):

similarly, (g o f)(x)= g(f(x)) so in g(x) put f(x) in place of x. try it.